I really need help.
n+2/15=n-4/25 they are fractions
Thanks
-Drea

Answer by Alan3354(21608)

n+2/15=n-4/25
--------------
Multiply by 75
5(n+2) = 3(n-4)
5n+10 = 3n-12
2n = -2
n = -1